OpenStudy (anonymous):

please help me.. i'm answering a take home exam in algebraic fractions.. i'm dying here.. :( i'm really really bad at math.. T~T

OpenStudy (jamesj):

If it's an exam, isn't it a violation of the honor code to ask us to help?

OpenStudy (anonymous):

it's just for reviewing purpose

OpenStudy (anonymous):

i need to practice and answer these equations

OpenStudy (jamesj):

Then post a question or two at the left.

OpenStudy (anonymous):

whets the ?

OpenStudy (anonymous):

\[[(x ^{2}-y ^{2})/x ^{3}-y ^{3}]\div[(x+y)/x]=?\]

OpenStudy (anonymous):

_____________

OpenStudy (anonymous):

please help me

hero (hero):

If you post it like this, it easier on the eyes: \[\frac{\frac{x^2-y^2}{x^3-y^3}}{\frac{x+y}{x}}\]